OpenStudy (anonymous):

Find the equation of the line that is parallel to the line 2x-3y=-4 containing the point (-5,3)

OpenStudy (karatechopper):

first you want to get it in slope intercept form

OpenStudy (anonymous):

ok

OpenStudy (karatechopper):

slope intercept form= y=mx+b

OpenStudy (karatechopper):

lets put this in slope intercept form! 2x-3y=-4 subtract 2x on both sides -3y=-2x-4 divide by negative 3 on both sides y-2/3x+4/3

OpenStudy (karatechopper):

mak that be y=2/3x+4/3

OpenStudy (saifoo.khan):

Ok, so we will start from solving that given line First of all, solve that line in the form of y=mx+c Where "m" is the slope. Note: Parallel line means same slope. 2x-3y=-4 Solve for y \[3y = 2x +4\]\[y = \frac23x + \frac43\] Comparing this with y=mx+c Slope is "m", so slope we get is 2/3. Now insert the point given and the slope in the formula, (i.e. slope and (-5,2) \[y-y_1 = m(x-x_1)\] \[y - 2 = \frac23 (x-(-5))\] Im sure u can solve now! :)

OpenStudy (karatechopper):

so now lets use a formula called point slope form. do you know the formula for that?

OpenStudy (anonymous):

how do you get y-2?

OpenStudy (saifoo.khan):

It's given in the question, Point (-5,2)

OpenStudy (anonymous):

its (-5,3)

OpenStudy (saifoo.khan):

Oh sorry. i overread. make it 3 please.

OpenStudy (anonymous):

\[y−2=2/3(x−(−5))\] y=2/3x +5.33

OpenStudy (saifoo.khan):

i prefer writing that in fractions

OpenStudy (anonymous):

is that the correct answer?

OpenStudy (saifoo.khan):

Nope.

OpenStudy (anonymous):

whats the answer?

OpenStudy (saifoo.khan):

\[y = \frac23 x+\frac{19}{3}\]

OpenStudy (anonymous):

its not 19/3. (2/3)*5+2= 5.33

OpenStudy (saifoo.khan):

hold on..

OpenStudy (anonymous):

ok?

OpenStudy (saifoo.khan):

lol, u made a mistake, (2/3)5 + 3

OpenStudy (anonymous):

hahah wow. Thanks for catching that!

OpenStudy (saifoo.khan):

;]